One Solution is 30% alcohol and another is 70% alcohol. Some of each solution is mixed to produce 8 liters of 40% solution. How many liters of the 70% solution should be used?
Answer by mananth(16946)

percentage---------------- quantity
Alcohol I70 ---------------- x liters
Alcohol II30 ---------------- 8 - x liters
Mixture 40.00% ---------------- 8

70x+30(8-x)=40*8

70x+240- 30x=320
70x-30x= 320-240
40x=80
/40
x=2 liters 70.00% Alcohol I
6 liters 30.00% Alcohol II